OpenStudy (anonymous):

the hypotenuse of a right triangle is 15 and that one leg is 5. What is the length of the other leg?

OpenStudy (asnaseer):

use the Pythagorus theorem - see here: http://www.mathsisfun.com/pythagoras.html

OpenStudy (poohbear):

You have to use the Pythagorean Theorem, which is A^2 + B^2 =C^2 . So the C would be the hypothenuse and B would be one of the legs. You put in the numbers. a^2 + 5^2 = 15^2. Then you square it. a^2 + 25 = 225. Then subtract 25 from 225. It becomes a^2 =200. Now, find the square root of 200, which is 14.14 and that rounds to 14. So, the other length would be 14.